OpenStudy (anonymous):

use Newtons method to find all roots of the equation; (x+1)^1/2-(x^2)+x

OpenStudy (goformit100):

@some_someone

OpenStudy (anonymous):

@goformit100 you have a greater Smart Score you should be able to help also :) 99 > 87

OpenStudy (anonymous):

\[\sqrt{x+1}+x^2+x=0\] is that what you are trying to solve?

OpenStudy (anonymous):

oh no, it is this \[\sqrt{x+1}-x^2+x=0\] right?

OpenStudy (anonymous):

first step is to make a guess it doesn't have to be a very good guess, but it is easier if it is good any ideas for a good starting point?

OpenStudy (anonymous):

guess not

OpenStudy (anonymous):

x=2, but I had to use the calculator to figure it out